--
JamieShiers - 28 Apr 2008
Week of 080428
Open Actions from last week:
Daily CCRC'08 Call details
To join the call, at 15.00 CET Tuesday to Friday inclusive (usually in CERN bat 28-R-006), do one of the following:
- Dial +41227676000 (Main) and enter access code 0119168, or
- To have the system call you, click here
CERN IT Status Board
Service Changes, Planned / Recent Interventions...
elogs
CCRC'08 elogs
Monday:
See the
weekly joint operations meeting minutes
Additional Material:
- CERN interventions:
- 29 April between 8:00 and 10:00 : lsf software upgarde and kernel upgrade on lsfmaster.
- 29 April at 9:30 There will be an upgrade of the CERN-PROD FTS services (FTS-T0-EXPORT, FTS-T1-IMPORT and FTS-T2-SERVICE) to the latest patch level at 09.30. There should be no user-visible downtime.
- 29 April upgrade of castor cms which should last approximately 3 hours.
Tuesday:
elog review: Nothing new.
Experiments round table:
ATLAS: This week is basically preparing for the CCRC'08 May run running the last functional tests of a new ddm release. This is being exercised exporting data at 40% of the full rate. There is currently a problem to NDGF in that other sites need the 64-bit ATLAS client to query for files there and it is no longer in the gLite distribution. We have obtained an rpm from NDGF and are testing it. In the long term NDGF will use an LFC so this problem will go away. Meanwhile simulation production for the June FDR continues. ATLAS now have the list of names of people authorised to raise operator alarms.
CMS: CERN Castor upgrade has just finished.
Sites round table:
CERN: there will be an emergency switch intervention at 09.30 tomorrow (Wednesday) lasting 5 minutes and cutting contact to several VO boxes (notably vocms06). See the CERN SLS pages from the IT home page.
RAL: is waiting to learn the experiment space token requests for the May run.
Core services (CERN) report:
DB services (CERN) report: There has been an intervention on the Oracle streams capture servers to install an Oracle patch to solve a problem needing an occassional manual restart of the servers. There was no downtime on the source databases. The replication of the LHCb LFC to FZK is occassionally losing contact (it is not known why only for this particular combination) - a call to Oracle has been opened. There will be a 3 days downtime at PIC during which we will desynchronise streams replication.
Monitoring / dashboard report:
Release update:
AOB:
Wednesday
elog review: Nothing new.
Experiments round table:
ALICE: will no longer support VO-boxes running under slc3. All Tier-1 are upgraded to slc4 but not all Tier-2. Sites can reuse their existing hardware.
ATLAS: From next week (start of CCRC'08 phase 2) NDGF will be running an LFC. Week 1 will be functional tests with data export to T1s at 40% of the full rate, week 2 will add T1 to T1 data movement then week 3 will start the full rate (even up to 120%) data export throughput tests. Note that Friday-Sunday each week are dedicated to detector commissioning.
Sites round table:
Core services (CERN) report:
DB services (CERN) report:
Monitoring / dashboard report:
Release update:
AOB: No meetings Thursday/Friday this week as CERN is closed.
Thursday
elog review:
Experiments round table:
Sites round table:
Core services (CERN) report:
DB services (CERN) report:
Monitoring / dashboard report:
Release update:
AOB: No meeting.
Friday
elog review:
Experiments round table:
Sites round table:
Core services (CERN) report:
DB services (CERN) report:
Monitoring / dashboard report:
Release update:
AOB: No meeting.